OpenStudy (anonymous):

solve each system by substitution be sure to give your solution as an ordered pair 3x+y=10 x-y=-6

OpenStudy (anonymous):

u multiply 3 * x?

OpenStudy (anonymous):

2 because 3 * 2 = 6 and 6 + 4 = 10

OpenStudy (anonymous):

so x = 2

OpenStudy (anonymous):

ok what about y

OpenStudy (anonymous):

where did u get the 2 from

OpenStudy (anonymous):

isnt x 1

OpenStudy (anonymous):

3x+y=10 x-y=-6 -------- 4x = 4 x = 1 x = 1 in x - y = -6 1 - y = -6 -y = -7 y = 7

OpenStudy (anonymous):

S={1,7}

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

Good

OpenStudy (anonymous):

can you help me work through one more though

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

x-y=1 x-3y=9

OpenStudy (anonymous):

make it 2x?

OpenStudy (anonymous):

yep, you can multiply one of them by -2

OpenStudy (anonymous):

so what would it look like

OpenStudy (anonymous):

-3 * -2?

OpenStudy (anonymous):

First, you can choose which equation do you want to multiply in order for you to exclude x. Let's try the first one. x-y=1 Now multiply by -2: -2(x) -2(y) = -2(1) Can you multiply it?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

you get -2x -2y = -2 did you understand why?

OpenStudy (anonymous):

u just multiply

OpenStudy (anonymous):

yup, you should have done that. now can you sum the multiplied equation with the second one? x will be 0 and you will find the value for y

OpenStudy (anonymous):

im lost now

OpenStudy (anonymous):

you have -2x -2y = -2 and x-3y=9. right?

OpenStudy (anonymous):

oh wait lol

OpenStudy (anonymous):

I thought it was another equation. Forget all of that. You want to solve x-y=1 x-3y=9 Just multiply the first or the second equation by -1.

OpenStudy (anonymous):

wait so forget bout everything u said b4 now

OpenStudy (anonymous):

Yep. Only consider my last post please

OpenStudy (anonymous):

x-y=1 (A) x-3y=9 (B) multiply (A) by -1: -x+y=-1 Now sum it with (B)

OpenStudy (anonymous):

okay well i figured it out its x = -3 and y = -4

OpenStudy (anonymous):

can you help me with this one x- y= 2 x+y=6

OpenStudy (anonymous):

Yup, that's correct. with this one you just have to do the same, actually it's easier because if you sum them you will find x

OpenStudy (anonymous):

2x?

OpenStudy (anonymous):

2x=8 now you have x=4 find y